OpenStudy (anonymous):

if y=mx+c, and x^2/a^2 + y^2/b^2 = 1, prove that c^2 - b^2 = (am)^2

OpenStudy (anonymous):

substitute for y and see what happens !

OpenStudy (anonymous):

i tried, some nonsense came, and it's too many variables turning my brain into mush. X_X

OpenStudy (anonymous):

lol

OpenStudy (sasogeek):

lol

OpenStudy (anonymous):

lol is RIGHT!

OpenStudy (anonymous):

A good one....

OpenStudy (anonymous):

!? :-/

OpenStudy (sasogeek):

I'm solving it but i think i made a mistake somewhere along the line... :/ i'll try again

OpenStudy (anonymous):

Problem is we have two c's?? One is distance from foci to centre =a^2-b^2 and the other hanging around in the line equation...

OpenStudy (anonymous):

Let's call the ellipse c C instead.

OpenStudy (anonymous):

Where's the op? Which c is what c? Not doing anything until u tell me.

OpenStudy (anonymous):

omg i tried so hard to solve this lol i think im missing something

OpenStudy (anonymous):

u c....:-)

OpenStudy (anonymous):

^LOL

OpenStudy (sasogeek):

I give up! darn it. anyone else still working on this? lemme know when u figure it out

OpenStudy (phi):

if y=mx+c, and x^2/a^2 + y^2/b^2 = 1, prove c^2 - b^2 = (am)^2 Consider m=1, c=2, a=2, b=1 See http://www.wolframalpha.com/input/?i=solve+y%3D+x%2B2%3B+%28x%2F2%29%5E2%2B%28y%2F1%29%5E2%3D1 then c^2-b^2 = 4-1= 3 (am)^2= (2*1)^2= 4 So unless we have an additional constraint, this is not true for the general case.

OpenStudy (anonymous):

Ah but which c is it?

OpenStudy (phi):

I only see 1 c in the given equations. Is there another one?

OpenStudy (sasogeek):

trying to solve and prove it, you get 2 c's, one is squared, the other isn't

OpenStudy (anonymous):

Some schools give focus distance as c^2 = a^2 -b^2..

OpenStudy (phi):

OK, I'll see if that makes any sense...

OpenStudy (anonymous):

I said to op I wouldn't do anything until she said which c was which...

OpenStudy (anonymous):

But mia since posting....

OpenStudy (phi):

If we assume c^2= a^2-b^2, then, in my test case, c^2- b^2= a^2-2b^2= 4-2=2 (am)^2= 4 No, it does not work!